# Pareto-Optimal Sampling and Resource Allocation for Timely Communication in Shared-Spectrum Low-Altitude Networks

### v1

@@ I Introduction
Low-altitude activities have grown significantly over the past decade, resulting in a surge in demand for communication links that support time-critical applications, such as real-time navigation, control, and surveillance [1]. The core requirement for these services is not merely reliability, but stringent data freshness. For instance, the value of a UAV’s field of view for mission monitoring is directly tied to its timeliness; a network that fails to deliver this information promptly renders it obsolete, severely compromising situational awareness and operational safety.
Most existing works on aerial communications focus on the throughput-reliability-delay trade-off [2, 3]. However, such approaches are insufficient to guarantee the stringent timeliness required by low-altitude networks. To address data freshness, some works employ age of information (AoI) to assess the importance of information and timing and to control sampling and transmission decisions, thereby balancing efficiency and timeliness [4, 5]. Taking Figure 1 as an example, status-aware sampling control can simultaneously reduce power consumption and RB allocation while guaranteeing timeliness. However, the high mobility of UAVs creates a time-varying topology with rapidly changing channel conditions for both communication and interference links. This highly dynamic environment presents a significant challenge in jointly optimizing sampling and resource allocation.
@@ II-A Predictive Channel Model
    advanced channel sensing techniques, such as radio maps and digital twins [12, 13], which provide a 3D representation of the wireless propagation environment and offer spatially resolved channel statistics between the UAV and ground BSs; and
@@ II-D Age-Aware Sampling Controller
The AoI serves not merely as a timeliness constraint; it also allows direct control of information generation [5, 15, 16]. Given the predictive communication model over the entire planning horizon, we aim to control the sampling timing of sensory data so as to achieve our goal with minimal payload.
@@ IV-B Problem Decomposition
Proof:
See Appendix C. ∎
As a result, instead of solving the Problem $\mathscr{P}2$ directly, we can solve $\mathscr{P}\text{2-1}$ and $\mathscr{P}\text{2-2}$ separately, with less variable coupling.
@@ V Graph-Based Algorithm for Optimal Control
This section proposes a framework to solve Problem $\mathscr{P}\text{2-1}$ and Problem efficiently $\mathscr{P}\text{2-2}$.
First, $\mathscr{P}\text{2-2}$ can be transformed into a convex problem by relaxing $a_{n}\left[k,t\right]\in[0,1]$ and introducing an auxiliary variable $\phi_{n}[k,t]=a_{n}\left[k,t\right]p_{n}\left[k,t\right]$. Consequently, $\mathscr{P}\text{2-2}$ can be solved with an optimality guarantee in $\mathcal{O}((t_{i+1}-t_{i})\theta\log_{2}((t_{i+1}-t_{i})\theta))$ complexity [8]. Thus, the remaining task is to solve the integer programming problem $\mathscr{P}\text{2-1}$.
@@ V-B2 Complexity Analysis
The complexity for all Pareto frontier searching is $\mathcal{O}(T\theta^{2}\bar{\tau}^{2}\log(\bar{\tau}\theta))$, where for each $\theta$, the complexity for the shortest path over a directed graph with nonnegative weights is $\mathcal{O}(T\bar{\tau}+T)=\mathcal{O}(T\bar{\tau})$, and the complexity for calculating all the weights $E^{*}(t_{i},t_{j})$ in $\mathscr{G}$ is $\mathcal{O}(T\theta\bar{\tau}^{2}\log\bar{\tau}\theta)$, which is proven in Appendix E.
@@ VI Simulation Results
We consider a UAV-based patrol system, where a UAV follows a circular trajectory to monitor a $200\times 200\,m^{2}$ area, as illustrated in Figure 5. The UAV operates at an altitude of $50$ m with a flight speed of 6 m/s. On the ground, $N$ BSs are deployed, with their positions randomly generated. The channels are realized according to (1). Similar to [8], the shape parameters $\kappa$ of the Gamma distribution of small-scale fading $\xi$ are set randomly in [1, 30], and the large-scale fading $g$ includes path loss and shadowing, where the path loss is generated by 3GPP Urban Micro (UMi) model [18] and the channel block state is generated by LOS probability model [19]. In contrast, the shadowing is modeled by a log-normal distribution.
@@
Figure 6 (a) illustrates the Pareto frontiers of all compared schemes. It can be observed that the non-predictive scheme (instantaneous rate) is significantly suboptimal, by more than 20 dB, compared with the other three predictive schemes, highlighting the importance of predictive information in strategy design. Moreover, although both the periodical and proposed status-aware schemes exploit predictive information, the latter achieves substantial performance improvement by adaptively controlling the sampling instants. For example, given an energy budget of $E=10\text{ dBm}$, the periodical sampling method requires approximately 300 RBs to complete the task. In contrast, the proposed status-aware method needs only about 50 RBs.

### v2

@@ I Introduction
Low-altitude activities have grown significantly over the past decade, resulting in a surge in demand for communication links that support time-critical applications, such as real-time navigation, control, and surveillance [1]. The core requirement for these services is not merely reliability, but stringent data freshness. For instance, the value of a unmanned aerial vehicle (UAV)’s field of view for mission monitoring is directly tied to its timeliness; a network that fails to deliver this information promptly renders it obsolete, severely compromising situational awareness and operational safety.
Most existing works on aerial communications focus on the throughput-reliability-delay trade-off [2, 3]. However, such approaches are insufficient to guarantee the stringent timeliness required by low-altitude networks. To address data freshness, some works employ age of information (AoI) and its variants to assess the importance of information and timing and to control sampling and transmission decisions, thereby balancing efficiency and timeliness [4, 5, 6, 7, 8, 9, 10]. Taking Figure 1 as an example, status-aware sampling control can simultaneously reduce power consumption and resource block (RB) allocation while guaranteeing timeliness. However, the high mobility of UAVs creates a time-varying topology with rapidly changing channel conditions for both communication and interference links. This highly dynamic environment presents a significant challenge in jointly optimizing sampling and resource allocation.
@@ II-A Predictive Channel Model
    advanced channel sensing techniques, such as radio maps and digital twins [17, 18, 19], which provide a 3D representation of the wireless propagation environment and offer spatially resolved channel statistics between the UAV and ground BSs; and
@@ II-D Age-Aware Sampling Controller
The AoI serves not merely as a timeliness constraint; it also allows direct control of information generation [10]. Given the predictive communication model over the entire planning horizon, we aim to control the sampling timing of sensory data so as to achieve our goal with minimal payload.
@@ IV-B Problem Decomposition
$\ \left\{p_{m}[k,t]\right\}\in\mathcal{P}(\mathcal{T}_{i}),\left\{a_{m}[k,t]\right\}\in\mathcal{A}(\mathcal{T}_{i})$
@@ V Graph-Based Algorithm for Optimal Control
This section proposes a framework to solve Problem $\mathscr{P}\text{2-1}$ and Problem $\mathscr{P}\text{2-2}$. First, $\mathscr{P}\text{2-2}$ can be transformed into a convex problem by relaxing $a_{n}\left[k,t\right]\in[0,1]$ and introducing an auxiliary variable $\phi_{n}[k,t]=a_{n}\left[k,t\right]p_{n}\left[k,t\right]$. Consequently, $\mathscr{P}\text{2-2}$ can be solved with an optimality guarantee in $\mathcal{O}((t_{i+1}-t_{i})\theta\log_{2}((t_{i+1}-t_{i})\theta))$ complexity [13]. Thus, the remaining task is to solve the integer programming problem $\mathscr{P}\text{2-1}$. The proofs can be found in [22].
@@ V-B2 Complexity Analysis
The complexity for all Pareto frontier searching is $\mathcal{O}(T\theta^{2}\bar{\tau}^{2}\log(\bar{\tau}\theta))$, where for each $\theta$, the complexity for the shortest path over a directed graph with nonnegative weights is $\mathcal{O}(T\bar{\tau}+T)=\mathcal{O}(T\bar{\tau})$, and the complexity for calculating all the weights $E^{*}(t_{i},t_{j})$ in $\mathscr{G}$ is $\mathcal{O}(T\theta\bar{\tau}^{2}\log\bar{\tau}\theta)$.
@@ VI Simulation Results
We consider a patrol system in which a UAV follows a circular trajectory to monitor a $200\times 200,\text{m}^{2}$ area, as illustrated in Figure 5. The UAV operates at an altitude of $50$ m with a flight speed of 6 m/s. On the ground, $N$ BSs are deployed, with their positions randomly generated. The channels are realized according to (1). Similar to [13], the shape parameters $\kappa$ of the Gamma distribution of small-scale fading $\xi$ are set randomly in [1, 30], and the large-scale fading $g$ includes path loss and shadowing, where the path loss is generated by 3GPP Urban Micro (UMi) model [23] and the channel block state is generated by LOS probability model [24]. In contrast, the shadowing is modeled by a log-normal distribution.
@@
Figure 6 (a) illustrates the Pareto frontiers of all compared schemes. It can be observed that the non-predictive scheme (instantaneous rate) is significantly suboptimal, by more than 20 dB, compared with the other three predictive schemes, highlighting the importance of predictive information in planning. Moreover, although both the periodical and proposed status-aware schemes exploit predictive information, the latter achieves substantial performance improvement by adaptively controlling the sampling instants. For example, given an energy budget of $E=10\text{ dBm}$, the periodical sampling method requires approximately 300 RBs to complete the task. In contrast, the proposed status-aware method needs only 50 RBs.
@@ Appendix C Proof of Proposition 4
$\ \left\{p_{m}[k,t]\right\}\in\mathcal{P}(\mathcal{T}_{i}),\left\{a_{m}[k,t]\right\}\in\mathcal{A}(\mathcal{T}_{i})$